OpenStudy (anonymous):

What are all the common factors of 36 and 56? A. 1, 2, 3, 4, 9 B. 1, 2 C. 1, 2, 4 D. 1, 2, 4, 8

OpenStudy (studygurl14):

36: what are the factors of 36? 1, 2, 3, 4, 6, 9, 12, 18, 36 Now, what are the factors of 56?

OpenStudy (jdoe0001):

\(\bf 36\implies {\color{blue}{ 2}}\cdot {\color{blue}{ 2}}\cdot 3\cdot 3\qquad 56\implies 2\cdot {\color{blue}{ 2}}\cdot {\color{blue}{ 2}}\cdot 7\) what do you think?

OpenStudy (anonymous):